# Dr.Mi-Bench: A Modular-integrated Benchmark for Scientific Deep Research Agent

### v1

@@ 7 Related Work
Early work on search agents, such as ReAct (yao2023react), established the paradigm of interleaving reasoning with tool use for multi-hop question answering (li2025search; jin2025search; xiong2025rag; wang-etal-2024-searching; xu2025collab; song2025r1; hao2025dynasearcher). While this line of work has been extended by commercial systems for more complex tasks (openai-blog; gemini-dr; perplexity-dr; grok-dr), they are either not designed for long-form synthesis (zhang2025web) or their evaluations are opaque and non-diagnostic. Similarly, existing benchmarks for DR are also limited; early systems like GAIA and BrowseComp focus on isolated retrieval or short-answer generation (mialon2023gaiabenchmarkgeneralai; wei2025browsecompsimplechallengingbenchmark; zhou2025browsecompzhbenchmarkingwebbrowsing; phan2025humanitysexam; futuresearch2025deepresearchbenchevaluating). Crucially, both lines of work fail to disentangle the distinct agent capabilities of planning, retrieval, and reasoning (see Appendix A.10 for more details).
@@ 8 Conclusion
To address critical gaps in the evaluation of DR agents, we introduce Dr.Mi-Bench, a human-annotated scientific benchmark, and Dr.Mi-Eval, a novel paradigm for the modular assessment of planning, retrieval, and reasoning. Our evaluation of state-of-the-art systems reveals a fragmented performance landscape where agents exhibit specialized strengths but also share critical weaknesses. By enabling the diagnostic analysis of such actionable failure modes, our work provides a crucial tool to guide the development of more reliable and trustworthy academic research assistants.
9 Ethics Statement
Our benchmark is constructed from publicly available, highly cited research papers across ten scientific disciplines. We ensure that all included materials are properly licensed or fall under fair use for academic research purposes. Nevertheless, several ethical considerations remain. First, as with any dataset derived from published literature, biases present in the source materials (e.g., disciplinary imbalance, geographic or linguistic concentration) may propagate into the benchmark. Second, while the benchmark focuses on scientific and technical domains, the reasoning outputs of DR agents could be misused to generate misleading or fabricated analyses if deployed irresponsibly. To mitigate these risks, we (i) release clear data statements describing the collection process, (ii) provide transparent evaluation protocols to discourage cherry-picking of results, and (iii) encourage responsible use of our benchmark for advancing trustworthy and reliable research assistants.
Appendix A Appendix
A.1 Deep Research Agent Module-wise Definitions
Definition 3 (Planner $\pi$)
@@ A.1 Deep Research Agent Module-wise Definitions
Definition 5 (Reasoner $\sigma$)

### v2

@@ 7 Related Work
Early work on search agents, such as ReAct (Yao et al., 2023), established the paradigm of interleaving reasoning with tool use for multi-hop question answering (Li et al., 2025a; Jin et al., 2025a; Xiong et al., 2025; Wang et al., 2024; Xu et al., 2025; Song et al., 2025; Hao et al., 2025). While this line of work has been extended by commercial systems for more complex tasks (OpenAI, 2025; Google, ; PerplexityAI, ; xAI, ), they are either not designed for long-form synthesis (Zhang et al., 2025b) or their evaluations are opaque and non-diagnostic. Similarly, existing benchmarks for DR are also limited; early systems like GAIA and BrowseComp focus on isolated retrieval or short-answer generation (Mialon et al., 2023; Wei et al., 2025; Zhou et al., 2025; Phan et al., 2025; FutureSearch et al., 2025). Crucially, both lines of work fail to disentangle the distinct agent capabilities of planning, retrieval, and reasoning (see Appendix B for more details).
@@ 8 Conclusion
To address critical gaps in the evaluation of DR agents, we introduce ADRA-Bank, a human-annotated academic benchmark, and ADRA-Eval, a novel paradigm for the modular assessment of planning, retrieval, and reasoning. Our evaluation of state-of-the-art systems reveals a fragmented performance landscape where agents exhibit specialized strengths but also share critical weaknesses. By enabling the diagnostic analysis of such actionable failure modes, our work provides a crucial tool to guide the development of more reliable and trustworthy academic research assistants.
Impact Statement
Our benchmark is constructed from publicly available, highly cited research papers across ten academic disciplines. We ensure that all included materials are properly licensed or fall under fair use for academic research purposes. Nevertheless, several ethical considerations remain. First, as with any dataset derived from published literature, biases present in the source materials (e.g., disciplinary imbalance, geographic or linguistic concentration) may propagate into the benchmark. Second, while the benchmark focuses on academic and technical domains, the reasoning outputs of DR agents could be misused to generate misleading or fabricated analyses if deployed irresponsibly. To mitigate these risks, we (i) release clear data statements describing the collection process, (ii) provide transparent evaluation protocols to discourage cherry-picking of results, and (iii) encourage responsible use of our benchmark for advancing trustworthy and reliable research assistants.
Appendix A Deep Research Agent Module-wise Definitions
Definition A.1 (Planner $\pi$).
@@ Appendix A Deep Research Agent Module-wise Definitions
Definition A.3 (Reasoner $\sigma$).